# AIReg-Bench: Benchmarking Language Models That Assess AI Regulation Compliance

### v1

@@ Abstract
As governments move to regulate AI, there is growing interest in using Large Language Models (LLMs) to assess whether or not an AI system complies with a given AI Regulation (AIR). However, there is presently no way to benchmark the performance of LLMs at this task. To fill this void, we introduce AIReg-Bench: the first benchmark dataset designed to test how well LLMs can assess compliance with the EU AI Act (AIA). We created this dataset through a two-step process: (1) by prompting an LLM with carefully structured instructions, we generated 120 technical documentation excerpts (samples), each depicting a fictional, albeit plausible, AI system — of the kind an AI provider might produce to demonstrate their compliance with AIR; (2) legal experts then reviewed and annotated each sample to indicate whether, and in what way, the AI system described therein violates specific Articles of the AIA. The resulting dataset, together with our evaluation of whether frontier LLMs can reproduce the experts’ compliance labels, provides a starting point to understand the opportunities and limitations of LLM-based AIR compliance assessment tools and establishes a benchmark against which subsequent LLMs can be compared. The dataset and evaluation code are available at https://github.com/camlsys/aireg-bench.
1 Introduction
@@
Across the world, AI Regulation (AIR) initiatives are either under development or have graduated the legislative process and gone into effect (Sloane & Wüllhorst, 2025; Chun et al., 2024; Alanoca et al., 2025). For both the regulators who enforce these regulations and the regulated parties who must comply with them, compliance assessments, whereby an AI system is evaluated for its compliance with respect to an AIR, play a pivotal role (Mökander et al., 2021; Ada Lovelace Institute, 2024; Anderljung et al., 2023; Raji et al., 2022; Reuel et al., 2024a). For example, the European Union’s AI Act (AIA) — dubbed “the world’s first comprehensive AI law” (European Parliament, 2024) — requires that providers of high-risk AI systems conduct such an assessment before putting their products on the market in the EU (EU, 2024, Art. 43).
@@
Despite their importance, however, certain AIR compliance assessments remain costly and time-consuming (Koh et al., 2024; Costanza-Chock et al., 2022; Sovrano et al., 2025). For example, some estimate that these assessments can take up to two-and-a-half days (European Commission, 2021) and cost EUR 7,500 for each AI system (Haataja & Bryson, 2021), accounting for up to 17% of the total expense of an AI project (Laurer et al., 2021). These high costs may contribute to a level of regulatory overhead that some have called unsustainable for AI providers and regulators alike (Laurer et al., 2021; Gikay, 2024; Reuel et al., 2024b; Koh et al., 2024; Molnar, 2024; Micklitz & Sartor, 2025) and, since it disproportionately affects small and medium-sized enterprises due to their lower resources (Stampernas & Lambrinoudakis, 2025), a potential hazard for fair competition (Martens, 2024; Gazendam et al., 2023; Wu & Liu, 2023; Guha et al., ; Iliasova et al., 2025).
@@
This may help explain why there is growing interest in, and experimentation with, using Large Language Models (LLMs) to perform (or, at least, streamline) AIR compliance assessments (Micklitz & Sartor, 2025; Li et al., 2025; Sovrano et al., 2025; Davvetas et al., 2025; Kővári et al., 2025; Makovec et al., 2024; Marino et al., 2024). And yet, there is still no standardized method for quantitatively evaluating and comparing the performance of LLMs at this particular task, creating uncertainty about the extent to which LLMs can be entrusted with it (Davvetas et al., 2025).
@@
To fill this void, we present AIReg-Bench: an open dataset for benchmarking the performance of LLMs at AIA compliance assessments. This dataset, which is available at https://github.com/camlsys/aireg-bench, consists of 120 technical documentation excerpts (i.e., details on system development and testing procedures) (Sovrano et al., 2025; Königstorfer & Thalmann, 2022). Each one provides information about a fictional, albeit plausible, AI system — specifically, a high-risk AI (HRAI) system under the AIA’s risk-based approach (EU, 2024, Art. 6). The samples in this dataset (i.e., the excerpts) are generated by an LLM-based technique, described in Section 2, allowing us to create diverse samples efficiently and at scale. As outlined in Section 3, each sample is then labeled by legal experts to indicate whether, and in what way, the system described therein violates specific Articles of the AIA.
@@
To showcase AIReg-Bench at work, we evaluated 10 frontier LLMs. Our findings indicate that some LLMs very closely approximate human expert judgments about the compliance (or lack thereof) of the excerpts in our dataset, such as Gemini 2.5 Pro (Comanici et al., 2025), which achieves a rank correlation of 0.856, as shown in Table 2.
@@
2 Sample Generation Pipeline
@@
When conducting a compliance assessment, regulators, consultants, and internal audit teams may draw upon many different kinds of input, including technical documentation, source code, and transcripts of staff interviews. However, in designing our benchmark, we optimized for simplicity and ease of use by focusing on a single type of input. Specifically, since AI compliance assessors have identified technical documentation as “the most important factor” in assessing whether an AI system complies with the governing regulations (Li & Goel, 2025) — something that we validated in our own interviews with AIR compliance experts (described in Appendix D) — we decided to make technical documentation (or excerpts thereof) the sole type of input to inform compliance assessments in our benchmark.
@@
The creation of our sample generation pipeline was motivated by two key bottlenecks: first, little to no AIA technical documentation of real AI systems is publicly available, perhaps due to the confidentiality or legal privilege surrounding such assessments (Guha et al., 2024); and second, paying experts to create them anew is prohibitively expensive (Pipitone & Alami, 2024). Therefore, inspired in part by Sovrano et al. (2025), who use an LLM to assist human drafting of technical documentation for AIA use cases, we designed a multi-stage pipeline, with gpt-4.1-mini (OpenAI, 2025) at its center, which generates plausible technical documentation excerpts efficiently and at scale. The stages of this pipeline are described in Section 2.1 and illustrated in Fig. 2.
@@
In devising this pipeline, we set forth several design criteria for the samples it generates. Many of these criteria, which are supplied in full in Appendix C, were written to ensure the benchmark remains manageably-scoped and easy to use. For example, we decided that our documentation should only depict HRAI systems within the scope of the AIA (EU, 2024, Art. 6). We also decided that each sample should be written only from the perspective of an AI system provider attempting to demonstrate compliance with a single article within the AIA (EU, 2024, Art. 2).
@@
Additionally, we labored to ensure our excerpts were realistic (i.e., representative of real-world technical documentation) and diverse (covering a range of AI systems with different intended uses and varying levels of compliance). As a concrete example, to give researchers control over the makeup and diversity of the distribution, we designed our pipeline to be steerable, allowing for the targeted generation of excerpts that are more or less likely to be compliant. The details of how this control was applied in AIReg-Bench are provided in Section 3, and further discussion of our design criteria and the rationale for these criteria can be found in Appendix C.
2.1 Stages of sample generation
  1. 1.
    First, gpt-4.1-mini is prompted to generate high-level overviews of AI systems, which fall into several use cases, such as road traffic control and credit scoring. By design, these use cases should be classified as high-risk under the AIA (EU, 2024, Art. 6(2); Ann. III).
  2. 2.
  3. 3.
The prompts used in each stage are included in Appendix E.
@@ 2.2 Validation of Plausibility
To validate the plausibility of the pipeline’s outputs, each excerpt in the AIReg-Bench dataset was reviewed by three legal experts (the same team of law school students, law graduates, and qualified lawyers who supplied the annotations for the dataset, described in Section 3). As described in the annotation instructions (Appendix F), these legal experts were asked to label each excerpt with the probability that it is plausible — i.e, that it is realistic, logically consistent, and reflects the type of technical documentation that a Fortune 500 Europe AI provider might realistically hand over to its compliance assessor or internal audit team. These labels help assess whether the process of using an LLM to generate technical documentation succeeded.
To be more specific, for each excerpt, annotators were asked to provide a score on a 1–5 Likert scale (Likert, 1932), where 1 indicated a very low probability of plausibility and 5 indicated a very high probability of plausibility (the exact phrasing of the Likert scale given in Appendix F). To accompany these quantitative scores, annotators were asked to provide a qualitative free text justification of the Likert score (i.e., a free text plausibility analysis). These scores and text entries are available in the AIReg-Bench GitHub repo.
The median plausibility score provided by annotators was 4 (i.e., high probability of plausibility).
3 Annotations and Dataset
Our second contribution involves using the sample generation pipeline that we created to generate a balanced distribution of samples, which legal experts then annotated in order to create the AIReg-Bench open benchmark dataset. This dataset is composed of 120 technical documentation excerpts that reflect 8 different use cases (intended uses) for AI systems and varying compliance profiles. A list of the use cases from the creation of AIReg-Bench is included as Appendix H and the prompt used to generate compliance profiles is included as Appendix E.2.
@@ 3 Annotations and Dataset
Table 1: AIReg-Bench dataset overview of human annotations for compliance on a Likert scale
@@ 3.1 Legal expert annotation of the excerpts
Generating annotations in the legal domain often demands specialized legal expertise reflecting deep subject-matter knowledge (Guha et al., 2024). Accordingly, multiple legal natural language processing and LLM benchmarks have leveraged legal expert annotators (including law school students and lawyers) (Wang et al., 2025b; Zheng et al., 2025; Östling et al., 2024; Wang et al., 2023; Hendrycks et al., 2021b; Leivaditi et al., 2020; Zhong et al., 2019; Duan et al., 2019; Wilson et al., 2016). Following this pattern, we used a team of six legal experts (law graduates, law students, or qualified lawyers) to review and annotate the technical documentation excerpts in AIReg-Bench. [^1] There are very few potential annotators with expertise in the EU AI Act who possess both the willingness and the capacity to carry out extended annotation tasks. For this reason, we broadened the eligibility criteria to include annotators with legal training more generally. Prior to annotation, these legal experts attended a training session, led by a legally-trained co-author, dedicated the AIA and its relevant articles. During annotation, labels were quality-checked by a law school graduate co-author.
In the end, each excerpt was reviewed by three legal experts. As described in the annotation instructions (Appendix F), for each excerpt, these annotators were asked to provide a score on a 1–5 Likert scale, where 1 indicated a very low probability of compliance with the relevant AIA article and 5 indicated a very high probability of compliance with that article (the exact phrasing of the Likert scale given in Appendix F). [^2] Annotators also mark whenever they found assigning a Likert score “difficult”. These markings can theoretically be used to segregate the more challenging samples. Notably, however, few annotators choose to use this demarcation in practice. To accompany these quantitative scores, annotators were asked to provide a qualitative free text justification of the Likert score (e.g., a free text compliance analysis). These justifications can, in theory, be semantically compared to the justifications produced by LLMs to provide another way to measure LLM compliance assessment capabilities.
4 Experiments
@@
In this evaluation, we prompted the LLMs to carry out the same task as the human annotators, supplying them with the identical documentation, system descriptions, and AIA articles that had been provided to the human annotators, along with instructions that were highly similar to those given to the human annotators (detailed in Appendix G). A key distinction, however, is that the annotators were free to consult external sources—such as websites or existing literature—whereas the LLMs were restricted to the materials explicitly supplied.
@@
o3 (OpenAI, 2025c) 0.723 0.809 -0.192 0.658
o3 mini (OpenAI, 2025b) 0.624 0.798 0.742 0.775
@@
Gemini 2.5 Pro (Comanici et al., 2025) 0.863 0.856 -0.225 0.458
Gemini 2.5 Flash (Comanici et al., 2025) 0.729 0.825 -0.108 0.625
Gemma 3 (Kamath et al., 2025) 0.696 0.757 0.258 0.692
@@
Despite prompts designed to mitigate sycophancy and acquiescence bias (Fanous et al., 2025), some models tended to assign higher compliance scores than human experts. Along this dimension, o3 mini and GPT-4o performed worst, with o3 mini strictly exceeding the median human expert score in 54.2% of excerpts, while only strictly falling below the median human expert’s score in 1.7% of excerpts (see Table 3).

### v3

@@ Abstract
As governments move to regulate AI, there is growing interest in using Large Language Models (LLMs) to assess whether or not an AI system complies with a given AI Regulation (AIR). However, there is presently no way to benchmark the performance of LLMs at this task. To fill this void, we introduce AIReg-Bench: the first open benchmark dataset designed to test how well LLMs can assess compliance with the EU AI Act (AIA). We created this dataset through a two-step process: (1) by prompting an LLM with carefully structured instructions, we generated 120 technical documentation excerpts (samples), each depicting a fictional, albeit plausible, AI system — of the kind an AI provider might produce to demonstrate their compliance with AIR; (2) legal experts then reviewed and annotated each sample to indicate whether, and in what way, the AI system described therein violates specific Articles of the AIA. The resulting dataset, together with our evaluation of whether frontier LLMs can reproduce the experts’ compliance labels, provides a starting point to understand the opportunities and limitations of LLM-based AIR compliance assessment tools and establishes a benchmark against which subsequent LLMs can be compared. The dataset and evaluation code are available at https://github.com/camlsys/aireg-bench.
Machine Learning, ICML
1 Introduction and Problem Statement
@@
Across the world, AI Regulation (AIR) initiatives are either under development or have graduated the legislative process and gone into effect (Sloane and Wüllhorst, 2025; Chun et al., 2024; Alanoca et al., 2025). For both the regulators who enforce these regulations and the regulated parties who must comply with them, compliance assessments, whereby an AI system is evaluated for its compliance with respect to an AIR, play a pivotal role (Mökander et al., 2021; Ada Lovelace Institute, 2024; Anderljung et al., 2023; Raji et al., 2022; Reuel et al., 2024a). For example, the European Union’s AI Act (AIA) — dubbed “the world’s first comprehensive AI law” (European Parliament, 2024) — requires that providers of high-risk AI systems conduct such an assessment before putting their products on the market in the EU (EU, 2024, Art. 43).
@@
Despite their importance, however, certain AIR compliance assessments remain costly and time-consuming (Koh et al., 2024; Costanza-Chock et al., 2022; Sovrano et al., 2025). For example, some estimate that these assessments can take up to two-and-a-half days (European Commission, 2021) and cost EUR 7,500 for each AI system (Haataja and Bryson, 2021), accounting for up to 17% of the total expense of an AI project (Laurer et al., 2021). These high costs may contribute to a level of regulatory overhead that some have called unsustainable for AI providers and regulators alike (Laurer et al., 2021; Gikay, 2024; Reuel et al., 2024b; Koh et al., 2024; Molnar, 2024; Micklitz and Sartor, 2025) and, since it disproportionately affects small and medium-sized enterprises due to their lower resources (Stampernas and Lambrinoudakis, 2025), a potential hazard for fair competition (Martens, 2024; Gazendam et al., 2023; Wu and Liu, 2023; Guha et al., ; Iliasova et al., 2025).
@@
This may help explain why there is growing interest in, and experimentation with, using Large Language Models (LLMs) to perform (or, at least, streamline) AIR compliance assessments (Micklitz and Sartor, 2025; Li et al., 2025; Sovrano et al., 2025; Davvetas et al., 2025; Kövári et al., 2025; Makovec et al., 2024; Marino et al., 2024; Falcarin et al., 2025; Videsjorden et al., 2026; Tran et al., 2025). And yet, there is still no standardized method for quantitatively evaluating and comparing the performance of LLMs at this particular task, creating uncertainty about the extent to which LLMs can be entrusted with it (Davvetas et al., 2025).
@@
To fill this void, we present AIReg-Bench: an open dataset for benchmarking the performance of LLMs at AIA compliance assessments. This dataset, which is available at https://github.com/camlsys/aireg-bench, consists of 120 technical documentation excerpts (i.e., details on system development and testing procedures) (Sovrano et al., 2025; Königstorfer and Thalmann, 2022). Each one provides information about a fictional, albeit plausible, AI system — specifically, a high-risk AI (HRAI) system under the AIA’s risk-based approach (EU, 2024, Art. 6). The samples in this dataset (i.e., the excerpts) are generated by an LLM-based technique, described in Section 2, allowing us to create diverse samples efficiently and at scale. As outlined in Section 3, each sample is then labeled by legal experts to indicate whether, and in what way, the system described therein violates specific Articles of the AIA.
@@
To showcase AIReg-Bench at work, we evaluated 10 frontier LLMs. Our findings indicate that some LLMs very closely approximate human expert judgments about the compliance (or lack thereof) of the excerpts in our dataset, such as Gemini 2.5 Pro (Comanici and others, 2025), which achieves a rank correlation of 0.856, as shown in Table 2.
@@
GPT-5GPT-4oo3o3 miniSonnet 4Gemini 2.5 ProGemini 2.5 FlashGemma 3Grok 4Grok 3 mini00.20.20.40.40.60.60.80.8110.850.850.770.770.720.720.620.620.770.770.860.860.730.730.70.70.830.830.730.73Cohen’s $\kappa$
@@
2 Sample Generation Pipeline (Solution Pt. I)
@@
When conducting a compliance assessment, regulators, consultants, and internal audit teams may draw upon many different kinds of input, including technical documentation, source code, and transcripts of staff interviews. However, in designing our benchmark, we optimized for simplicity and ease of use by focusing on a single type of input. Specifically, since AI compliance assessors have identified technical documentation as “the most important factor” in assessing whether an AI system complies with the governing regulations (Li and Goel, 2025) — something that we validated in our own interviews with AIR compliance experts (described in Appendix C) — we decided to make technical documentation (or excerpts thereof) the sole type of input to inform compliance assessments in our benchmark.
@@
The creation of our sample generation pipeline was motivated by two key bottlenecks: first, little to no AIA technical documentation of real AI systems is publicly available, perhaps due to the confidentiality or legal privilege surrounding such assessments (Guha et al., 2024); and second, paying experts to create them anew is prohibitively expensive (Pipitone and Alami, 2024). Therefore, inspired in part by Sovrano et al. (2025), who use an LLM to assist human drafting of technical documentation for AIA use cases, we designed a multi-stage pipeline, with gpt-4.1-mini (OpenAI, 2025c) at its center, which generates plausible technical documentation excerpts efficiently and at scale. The stages of this pipeline are described in Section 2.1 and illustrated in Fig. 2. Although we also experimented with o3-mini (OpenAI, 2025d), 4o-mini (OpenAI, 2024), and gpt-5 (OpenAI, 2025a), gpt-4.1-mini was ultimately chosen for the pipeline due to its price point and the satisfactory outputs produced during our experimentation phase.
@@
In devising this pipeline, we set forth several design criteria for the samples it generates. Many of these criteria, which are supplied in full in Appendix B, were written to ensure the benchmark remains manageably-scoped and easy to use. For example, we decided that our documentation should only depict HRAI systems within the scope of the AIA (EU, 2024, Art. 6). Although we could have chosen any subset of AIA requirements to treat as a proof of concept for this benchmarking effort, we felt the AIA’s HRAI requirements were an especially worthwhile subject because many regard them as “the most important part of the AIA” (Araszkiewicz et al., 2022). We also decided that each sample should be written only from the perspective of an AI system provider attempting to demonstrate compliance with a single article within the AIA (EU, 2024, Art. 2).
@@
Additionally, we labored to ensure our excerpts were realistic (i.e., representative of real-world technical documentation) and diverse (covering a range of AI systems with different intended uses and varying levels of compliance). As a concrete example, to give researchers control over the makeup and diversity of the distribution, we designed our pipeline to be steerable, allowing for the targeted generation of excerpts that are more or less likely to be compliant. The details of how this control was applied in AIReg-Bench are provided in Section 3, and further discussion of our design criteria and the rationale for these criteria can be found in Appendix B.
@@
2.1 Stages of sample generation
  1. 1.
    First, gpt-4.1-mini is prompted to generate high-level overviews of AI systems, which fall into several use cases, such as road traffic control and credit scoring. By design, these use cases should be classified as high-risk under the AIA (EU, 2024, Art. 6.2; Ann. III).
  2. 2.
  3. 3.
The prompts used in each stage are included in Appx. D.
@@ 2.2 Validation of Plausibility
To validate the plausibility of the pipeline’s outputs, each excerpt in the AIReg-Bench dataset was reviewed by three legal experts (the same team of law school students, law graduates, and qualified lawyers, all regulation specialists, who supplied the annotations for the dataset, described in Section 3). As described in the annotation instructions (Appendix E), these legal experts were asked to label each excerpt with the probability that it is plausible — i.e, that it is realistic, logically consistent, and reflects the type of technical documentation that a Fortune 500 Europe AI provider might realistically hand over to its compliance assessor or internal audit team. These labels help assess whether the process of using an LLM to generate technical documentation succeeded.
To be more specific, for each excerpt, annotators were asked to provide a score on a 1–5 Likert scale (Likert, 1932), where 1 indicated a very low probability of plausibility and 5 indicated a very high probability of plausibility (the exact phrasing of the Likert scale given in Appendix E). To accompany these quantitative scores, annotators were asked to provide a qualitative free text justification of the Likert score (i.e., a free text plausibility analysis). These scores and text entries are available in the AIReg-Bench GitHub repo. Notably, the median plausibility score provided by annotators was 4 (i.e., high probability of plausibility).
3 Annotations and Dataset (Solution Pt. II)
Our second contribution involves using the sample generation pipeline that we created to generate a balanced distribution of samples, which legal experts then annotated in order to create the AIReg-Bench open benchmark dataset. This dataset is composed of 120 technical documentation excerpts that reflect 8 different use cases (intended uses) for AI systems and varying compliance profiles. A list of the use cases from the creation of AIReg-Bench is included as Appendix G and the prompt used to generate compliance profiles is included as Appendix D.2.
@@
Table 1: AIReg-Bench dataset overview of human annotations for compliance on a Likert scale. The uses correspond to, in order: traffic safety, gas delivery, education, exam proctoring, job hiring, job termination, emergency dispatch, and credit scoring.
@@ 3.1 Legal expert annotation of the excerpts
Generating annotations in the legal domain often demands specialized legal expertise reflecting deep subject-matter knowledge (Guha et al., 2024). Accordingly, multiple legal natural language processing and LLM benchmarks have leveraged legal expert annotators (including law school students and lawyers) (Wang et al., 2025b; Zheng et al., 2025; östling2024cambridgelawcorpusdataset; Wang et al., 2023; Shen et al., 2022; Hendrycks et al., 2021b; Leivaditi et al., 2020; Zhong et al., 2019; Duan et al., 2019; Wilson et al., 2016). Following this pattern, we used a team of six legal experts (law graduates, law students, or qualified lawyers) to review and annotate the technical documentation excerpts in AIReg-Bench. [^1] There are very few potential annotators with expertise in the EU AI Act who possess both the willingness and the capacity to carry out extended annotation tasks. For this reason, we broadened the eligibility criteria to include annotators with legal training more generally — although, ultimately, our entire team listed regulation as one of their specializations (see Appendix H). As detailed in Appendix H, each of these annotators listed regulation as one of their specializations. Prior to annotation, these legal experts attended a training session, led by a legally-trained co-author, dedicated to the AIA and its relevant articles. During annotation, labels were quality-checked by a law school graduate co-author.
In the end, each excerpt was reviewed by three legal experts. As described in the annotation instructions (Appendix E), for each excerpt, these annotators were asked to provide a score on a 1–5 Likert scale, where 1 indicated a very low probability of compliance with the relevant AIA article and 5 indicated a very high probability of compliance with that article (the exact phrasing of the Likert scale given in Appendix E). [^2] Annotators also mark whenever they found assigning a Likert score “difficult”. These markings can theoretically be used to segregate the more challenging samples. Notably, however, few annotators choose to use this demarcation in practice. To accompany these quantitative scores, annotators were asked to provide a qualitative free text justification of the Likert score (e.g., a free text compliance analysis). These justifications can, in theory, be semantically compared to the justifications produced by LLMs to provide another way to measure LLM compliance assessment capabilities.
Appendix I contains a select annotation example. Appendix J contains an analysis of recurring themes in these annotations. The annotators’ inter-rater reliability, as measured with Krippendorff’s alpha coefficient (Krippendorff, 2018), was 0.651. This suggests a moderate level of agreement, though too low for drawing strong conclusions, likely reflecting the subjectivity of legal judgements.
Variance is heavily structured by two annotators with opposing biases (-0.917 and +0.600). Removing these annotators increases Krippendorff Alpha by +0.1343 to 0.786. That said, for the analyses in this paper, we retained all annotations to avoid post hoc exclusion.
The highest average disagreement occurred around Article 10 and Article 15, with mean standard deviations of 0.638 and 0.612 respectively. The most severe disagreement involved the intended use of Credit Scoring (Use 8) in light of Article 9, whose scores (1, 5, 5) resulted in the dataset’s highest standard deviation (1.89).
4 Experiments (Proving Our Solutions Address the Problem)
@@
In this evaluation, we prompted the LLMs to carry out the same task as the human annotators, supplying them with the identical documentation, system descriptions, and AIA articles that had been provided to the human annotators, along with instructions that were highly similar to those given to the human annotators (detailed in Appendix F). A key distinction, however, is that the annotators were free to consult external sources—such as websites or existing literature—whereas the LLMs were restricted to the materials explicitly supplied.
@@
o3 (OpenAI, 2025b) 0.723 0.809 -0.192 0.658
o3 mini (OpenAI, 2025d) 0.624 0.798 0.742 0.775
@@
Gemini 2.5 Pro (Comanici and others, 2025) 0.863 0.856 -0.225 0.458
Gemini 2.5 Flash (Comanici and others, 2025) 0.729 0.825 -0.108 0.625
Gemma 3 (Kamath and others, 2025) 0.696 0.757 0.258 0.692
@@
Despite prompts designed to mitigate sycophancy and acquiescence bias (Fanous et al., 2025), some models tended to assign higher compliance scores than human experts. Along this dimension, o3 mini and GPT-4o performed worst, with o3 mini strictly exceeding the median human expert score in 54.2% of excerpts, while only strictly falling below the median human expert’s score in 1.7% of excerpts (Table 3).